Grade 10Physical Chemistry

a sample of drinking water was found to be severely contaminated with chloroform CHCl 3 supposed to be a carcinogen the level of contamination was 15ppm by mass
  1. express this in % mass
  2. determine the molality of chloroform in the water sample

Aiden Mathew Denny
7 Years agoGrade 10

2 Answers

Arun
7 Years ago
Dear student
 
 
1) 15 ppm means 15 parts in 1 millions (10^6)
and percentage always for 100
 
hence 15 ppm = (15/10^6 )*100
= 15*10^-4
= 1.5*10^-3
2) molar mass of chloroform = 118.5g/ mol
 
from percentage by mass 100 gram of sample has 1.5*10^-3 g of chloroform
then 1000 gram of sample has = 1.5*10^-3 g * 1000 gram/ 100 gram
= 1.5*10^-2 g
 
molality = moles of solute / mass of solvent in Kg
 
moles of solute = mass of solute / molar mass
= 1.5*10^-2/ 118.5 
=1.266*10^-4 moles
 
molality =  1.266*10^-4 moles/ 1 Kg =1.266*10^-4 m
 
Regards
Arun
Soumendu Majumdar
7 Years ago
Dear Aiden,
PPM refers to parts per million or (mass of chloroform) x 10^{6} divided by the total mass--Equation 1
Mass percentage refers to the mass of chloroform x 10^{2} divided by the total mass--Equation 2
So dividing Equation 2 by Equation 1 we get
Mass percentage =  15\times 10^{2}/ 10^6 = 15\times 10^{-4
Now molality is defined as the ratio of the number of moles of solute divided by the mass of solvent in kg
The molecular mass of chloroform or mass of 1 mole of chloroform is 119.38 g/mol (you may round it off as 119.5 or 120 g/mol as required)
Now out of 100g of the solution, the mass of chloroform is 15\times 10^{-4 g
So the mass of solvent = 100-15\times 10^{-4 g = 99.9985 g (You may round it off as 100 g)
= 0.0999985 kg (can be rounded off to 0.1 kg)
Now the number of moles of chloroform = 15\times 10^{-4}/119.38
So molality = 15\times 10^{-4}/119.38 \times 0.0999985 mol/kg
1.2565\times 10^{-4} mol/kg
Hope this helps!
regards,
Soumendu